# Multiscale Training of Convolutional Neural Networks

### v1

@@ 2.2 Multiscale Analysis of Convolutional Neural Networks
Lemma 1 (Convergence of standard convolution kernels.).
@@
$\|{\bf g}^{2h}-{\bf g}^{h}\|={\cal O}(h).$
As can be shown in the proof (see Appendix B), the convergence of the gradient depends on the amount of noise in the data. However, as we now show in Example 1, for noisy problems, this method may not always yield the desired outcome.
@@
where ${\bf R}_{h}^{2^{i}h}$ is a linear interpolation operator that takes the signal from mesh size $h$ to mesh size $2^{i}h$. We compute the difference between the gradient of the function on each level and its norm
$\delta g=\|{\bm{\nabla}\,}_{{\bm{\theta}}}\ell_{2^{i}h}-{\bm{\nabla}\,}_{{\bm{\theta}}}\ell_{2^{i+1}h}\|^{2}.$ (16)
We conduct two experiments. In the first, we use smooth inputs, and in the second, we add Gaussian noise $N(0,1)$ to the inputs. We compute the values of $\delta g$, and report it in Table 1. The experiment demonstrates that for noisy inputs, the gradient with respect to the convolution parameters grows in tandem with the mesh. Thus, using convolution in its standard form will not be useful in reducing the error using the telescopic sum in Equation 6 as described above.
@@
Noisy input $9.3\times 10^{-3}$ $7.1\times 10^{-3}$ $4.9\times 10^{-3}$ $2.8\times 10^{-3}$ $2.0\times 10^{-3}$ $1.7\times 10^{-3}$
Smooth input $1.2\times 10^{-4}$ $2.5\times 10^{-4}$ $5.1\times 10^{-4}$ $1.0\times 10^{-3}$ $1.9\times 10^{-3}$ $3.7\times 10^{-3}$
Table 1: The discrepancy between two consecutive gradients for the problem in Equation 14. As the mesh resolution increases, the discrepancy reduces for smooth data, while it increases for noisy data.
The analysis suggests that for smooth signals, standard CNNs can be integrated into Multilevel Monte-Carlo methods, which form the basis of our Multiscale-SGD. However, in many cases, inputs may be noisy or contain high frequencies, leading to large or unbounded gradients across mesh resolutions. This explains the lack of convergence of standard CNNs in Example 1. High-frequency inputs can hinder learning in multiscale frameworks like Multiscale-SGD. To address this, we propose a new convolution type in Section 3 that theoretically alleviates this issue.
@@ 3.1 Mesh-Free Convolutions
We introduce Mesh-Free Convolutions (MFCs), which define convolutions in the functional space independently of any mesh, then discretize them on a given mesh. Unlike standard convolution kernels tied to specific meshes, MFCs converge to a finite value as the mesh refines, overcoming the limitations of standard kernels discussed in Section 2.2. One way to achieve this goal is to define the convolution in functional space directly by an integral, that is
@@
where the kernel function ${\cal K}_{{\bm{\xi}}}$ is, parametrized by learnable parameters ${\bm{\xi}}$. This general convolution can be clearly defined on any mesh and is basically mesh-free. The question remains, how to parametrize the kernel. Clearly, there are many options for this choice. A too restrictive approach leads to limited expressivity of the convolution. Our goal and intuition is to generate kernels that are commonly obtained by trained CNNs. Such filters are rarely random but contain some structure (Goodfellow et al., 2016). To this end, we use the interplay between the solution of partial differential equations (PDEs) and convolutions (Evans, 1998). In particular, the solution of parabolic PDEs (that is, the heat equation) is a kernel that exhibits the behavior that is often required from CNN filters.
Let $u$ be input functions (that is, $u$ can have more than a single channel, or so-called vector function) and let $v={\cal C}({\bm{\xi}})u$ be the application of the mesh-free convolution is parameterized by ${\bm{\xi}}$. We obtain a mesh-independent convolution by composing two processes. The first is a trainable matrix ${\bf R}\in\mathbb{R}^{d\times c}$ which is a $1\times 1$ convolution, that embeds $u$ into a subspace ${\tilde{v}}$, yielding the initial condition of the PDE defined in Equation 18. The second is the solution of the parabolic PDE of the form below, which maps the initial condition $\tilde{v}(0,{\bf x})$ into $\tilde{v}(\tau,{\bf x})$
$\tilde{v}_{t}={\cal L}({\bm{\alpha}})\tilde{v},\quad\tilde{v}(t=0)={\bf R}u,\quad t\in[0,\tau],\quad\quad v=(\gamma_{0}+[\gamma_{x},\gamma_{y}]^{\top}{\bm{\nabla}\,})\tilde{v}({\bf x},\tau).$ (18)
Finally, the output $v({\bf x})$ is obtained by taking a directional derivative of $\tilde{v}(\tau,{\bf x})$ in a direction ${\bm{\gamma}}$. All the processes above are encompassed by the mesh-free convolution operator ${\cal C}({\bm{\xi}})$. The operator ${\cal L}({\bm{\alpha}})={\rm diag}\left({\cal L}({\bm{\alpha}}_{1}),\ldots,{\cal L}({\bm{\alpha}}_{d})\right)$ is an Elliptic Differential Operator, that we define as
@@
The trainable parameters are ${\bm{\xi}}=[\tau,{\bm{\alpha}},{\bm{\gamma}}]\in\mathbb{R}^{7}$, where $\tau$ is a scalar, ${\bm{\gamma}}=[\gamma_{0},\gamma_{x},\gamma_{y}]$, and ${\bm{\alpha}}=[\alpha_{xx},\alpha_{xy},\alpha_{yy}]$. We enforce $\alpha_{xx},\alpha_{yy}>0$ and $\alpha_{xx}\alpha_{yy}-\alpha_{xy}^{2}>0$, ensuring ${\cal L}$ is elliptic and invertible (Trottenberg et al., 2001). Integrating the PDE over time $[0,\tau]$ and differentiating in the direction $[\gamma_{x},\gamma_{y}]$ effectively applies the convolution to the input $u$. The parameter $\tau$ acts as a scaling factor: increasing it widens and smooths the kernel, while as $\tau\to 0$, the kernel becomes more localized. To visualize our convolution, we sample random parameters ${\bm{\xi}}$ and plot the resulting MFC filters in Figure 3. By varying ${\bm{\xi}}$, we generate a range of convolutional kernels, from local to global, with different orientations. Here, $\alpha$ controls orientation, $\gamma$ sets the lobe direction, and $\tau$ determines the kernel width. The figure demonstrates that our choice of convolutions yields trainable filters to the ones that are typically obtained by CNN (Goodfellow et al., 2016). These filters can change their field of view by changing $t$, change their orientation by changing ${\bm{\xi}}$, and change their side lobes’ strength and direction by changing ${\bm{\gamma}}$.
@@
Figure 3: An illustration of Mesh-Free Convolutions (MFCs) is provided to showcase the effect of parameter variations. First row: The parameters are set as $\alpha_{xx}=1$, $\alpha_{yy}=3$, while $\alpha_{xy}$ varies from 0 to 1, and ${\bm{\gamma}}_{x}$ changes from 0 to 1. Second row: The parameters are set as ${\bm{\gamma}}_{x}={\bm{\gamma}}_{y}=0$, ${\bm{\gamma}}_{0}=1$, with $\alpha_{xy}$ varying from 0 to 1, and $t$ changing from 0 to 1. These visualizations highlight the flexibility and expressiveness of MFCs in adapting to different parameter configurations.
@@
Computing MFCs on a Discretized Mesh. Our MFCs are defined by the solution of Equation 18 and are continuous functions. Here, we discuss their computation on a discretized mesh for application to images. For a rectangular domain $\Omega$, we use periodic boundary conditions and the Fast Fourier Transform (FFT) to compute the convolution. Let ${\bf u}^{h}$ be a discrete tensor. First, we compute its Fourier transform, $\widehat{\bf u}^{h}={{\bf F}}{\bf u}^{h}$, and use a frequency mesh to obtain the 2D frequency grid ${\bf k}_{x}$ and ${\bf k}_{y}$. The discrete symbol ${\cal L}({\bf k}_{x},{\bf k}_{y})$ is a 3D tensor operating on each channel, similar to standard 2D filters. We then compute a point-wise product and use the inverse FFT to return to image space. Interestingly, given some fixed mesh of size $h$ and a regular 2D convolution ${\bf C}$ it is always possible to build an MFC, ${\bf C}_{{\bm{\xi}}}$, such that they are equivalent when evaluated on that mesh. The proof is given in Appendix B, and it implies that MFCs have at least the same expressiveness as standard convolutions.
@@ 3.2 Fourier Analysis of Mesh-Free Convolutions
Following the analysis path shown in Section 2.2, consider the linear loss obtained by these MFCs when discretizing the functions $u$ and $y$. In the case of functions, such as with our MFCs, the inner product in Equation 14 is replaced by an integral that reads:
@@
We study the behavior of our MFC, denoted by ${\cal C}({\bm{\xi}})$, on a function $u(x)$, by computing its symbol. Because the heat equation is linear, the symbol of ${\cal C}({\bm{\xi}})$ is given by (see Trottenberg et al. (2001)):
${\cal L}(k_{x},k_{y})=\left(\gamma_{0}+ik_{x}\gamma_{x}+ik_{y}\gamma_{y}\right)\exp\left(-\tau(\alpha_{xx}k_{x}^{2}+2\alpha_{xy}k_{x}k_{y}+\alpha_{yy}k_{y}^{2})\right).$ (21)
@@
Lemma 2 (MFCs yield consistent gradients independently of mesh resolution.).
@@
The property described in Lemma 2 carries forward to any discrete space, i.e., chosen mesh. Namely, by discretizing our MFCs on a uniform mesh with mesh-size $h$ we obtain its discrete analog:
@@
$\ell^{h}({\bm{\xi}})={\frac{1}{n}}({\bf F}{\bf y}^{h})^{\top}\big{(}{\cal L}(k_{x},k_{y}){\bf F}{\bf u}^{h}\big{)}={\frac{1}{n}}\sum_{k_{x},k_{y}}{\cal L}(k_{x},k_{y})\odot\widehat{\bf y}^{h}(k_{x},k_{y})\odot\widehat{\bf u}^{h}(k_{x},k_{y}).$ (23)
As both $\widehat{\bf y}^{h}$ and $\widehat{\bf u}^{h}$ are discrete Fourier transforms of the functions $y$ and $u$, they converge to their continuous counterparts as $h\rightarrow 0$. Thus, $\ell^{h}$ converges to $\ell^{*}$ as $h\rightarrow 0$, leading to a key result: unlike standard convolutions, the gradients of our MFCs converge to a finite value as the mesh refines, making them suitable for multiscale techniques like Multiscale-SGD, summarized in Lemma 3.
Lemma 3 (MFCs gradients converge as the mesh resolution refines.).
Let $\ell^{h}({\bm{\xi}})$ be define by Equation 22. Then at the limit,
@@
Noisy input $1.6\times 10^{-4}$ $3.2\times 10^{-4}$ $6.1\times 10^{-4}$ $1.3\times 10^{-3}$ $2.9\times 10^{-3}$ $5.9\times 10^{-3}$
Smooth input $1.1\times 10^{-4}$ $2.1\times 10^{-4}$ $4.2\times 10^{-4}$ $7.8\times 10^{-4}$ $1.5\times 10^{-3}$ $3.1\times 10^{-3}$
Table 2: The discrepancy between two consecutive gradients for the example problem Equation 14 using mesh-free convolutions. The discrepancy is reduced for both noisy and smooth data.
@@ 4 The Full-Multiscale Algorithm
where $C$ is a constant, ${\bm{\theta}}_{*}^{h}$ is the parameter that optimizes the expectation of the loss, and $C_{0}$ is a constant that depends on the initial error at ${\bm{\theta}}_{0}$. Let ${\bm{\theta}}_{*}^{H}$ and ${\bm{\theta}}_{*}^{h}$ be the parameters that minimize the expectation of the loss for meshes with resolution $H>h$ and assume that $\|{\bm{\theta}}_{*}^{H}-{\bm{\theta}}_{*}^{h}\|\leq\gamma H$, where $\gamma$ is a constant. This assumption is justified if the loss converges to a finite value as $h\rightarrow 0$. During the Full-Multiscale-SGD iterations (Algorithm 2), we solve the problem on the coarse mesh $H$ to initialize the fine mesh $h$. Thus, after $k$ steps of SGD on the fine mesh, we can bound the error by:
$\left|\left[{\mathbb{E}}{\ell}({\bm{\theta}}_{k})\right]-{\mathbb{E}}\left[{\ell}({\bm{\theta}}_{*}^{h})\right]\right|\leq\gamma H{\frac{C}{k}}.$ (25)
Requiring that the error is smaller than some $\epsilon$, renders a bounded number of required iterations $k\approx\gamma C{\frac{H}{\epsilon}}.$ Since in Algorithm 2, $H=2^{j}h$, the number of iterations for a fixed error $\epsilon$ is halved at each level, the iterations on the finest mesh are a fraction of those on the coarsest mesh. As shown in Table 3, multiscale speeds up training by a factor of $\sim$160 compared to standard single-scale training.

### v2

@@ 2.2 Multiscale analysis of convolutional neural networks
We now analyze under which conditions Multiscale-SGD can be applied for neural network optimization without compromising on its efficiency.
@@
Lemma 2.1 (Convergence of standard convolution kernels).
@@
$\|{\bf g}^{2h}-{\bf g}^{h}\|\sim{\cal O}(h).$
As shown in the proof (see Appendix B), the convergence of the gradients depends on the amount of noise in the data. However, as we now show in Example 1, for noisy problems, this method may not always yield the desired outcome.
@@
where ${\bf R}_{h}^{2^{i}h}$ is a linear interpolation operator that takes the signal from mesh size $h$ to mesh size $2^{i}h$. We compute the difference between the gradient of the function on each level and its norm, $\delta g=\|{\bm{\nabla}\,}_{{\bm{\theta}}}\ell_{2^{i}h}-{\bm{\nabla}\,}_{{\bm{\theta}}}\ell_{2^{i+1}h}\|^{2}$.
We conduct two experiments: in the first, we use smooth inputs, and in the second, we use smooth inputs with added Gaussian noise $\mathcal{N}(0,1)$. We compute the values of $\delta g$, and report it in Table 1. The experiment demonstrates that for noisy inputs, the gradient with respect to the convolution parameters grows in tandem with the mesh. Thus, using convolution in its standard form will not be useful in reducing the error using the telescopic sum in Section 2.1 as described above.
Table 1: The discrepancy between two consecutive gradients at different resolutions for the loss in Equation 13 for standard convolutions. As the mesh resolution increases, the discrepancy reduces for smooth data, while it increases for noisy data.
Computed loss on different resolutions (×$\times 10^{-3}$)
@@
Noisy input 9.3 7.1 4.9 2.8 2.0 1.7
Smooth input 0.12 0.25 0.51 1.0 1.9 3.7
The analysis suggests that for smooth signals, standard CNNs can be integrated into Multilevel Monte-Carlo methods, which form the basis of our Multiscale-SGD. However, in many cases, inputs may be noisy or contain high frequencies, leading to large or unbounded gradients across mesh resolutions. This explains the lack of convergence of standard CNNs in Example 1. High-frequency inputs can hinder learning in multiscale frameworks like Multiscale-SGD. To address this, we propose a new convolution type in Section 3 that theoretically alleviates this issue when dealing with very noisy data.
@@ 3.1 Mesh-Free Convolutions
We introduce Mesh-Free Convolutions (MFCs), which define convolutions in the functional space independently of any mesh, then discretize them on a given mesh obtaining consistency. Unlike standard convolution kernels tied to specific meshes, MFCs converge to a finite value as the mesh refines, overcoming the limitations of standard kernels discussed in Section 2.2. One way to achieve this goal is to define the convolution in functional space directly by an integral,
@@
where the kernel function ${\cal K}_{{\bm{\xi}}}$ is, parametrized by learnable parameters ${\bm{\xi}}$. This general convolution can be defined on any mesh and is basically mesh-free. The question remains: how to parametrize the kernel? Clearly, there are many options for this choice. A too-restrictive approach leads to limited expressivity of the convolution. Our goal and intuition is to generate kernels that are commonly obtained by trained CNNs. Such filters are rarely random but contain some structure (Goodfellow et al., 2016). To this end, we use the interplay between the solution of partial differential equations (PDEs) and convolutions (Evans, 1998). In particular, the solution of parabolic PDEs (that is, the heat equation) is a kernel that exhibits the behavior that is often required from CNN filters.
Let $u$ be input functions (that is, $u$ can have more than a single channel, or so-called vector function) and let $v={\cal C}({\bm{\xi}})u$ be the application of the mesh-free convolution is parameterized by ${\bm{\xi}}$. We obtain a mesh-independent convolution by composing two processes. The first is a trainable matrix ${\bf R}\in\mathbb{R}^{d\times c}$ which is a $1\times 1$ convolution, that embeds $u$ into a subspace ${\tilde{v}}$, yielding the initial condition of the PDE defined in Section 3.1. The second is the solution of the parabolic PDE of the form below, which maps the initial condition $\tilde{v}(0,{\bf x})$ into $\tilde{v}(\tau,{\bf x})$
$\tilde{v}_{t}={\cal L}({\bm{\alpha}})\tilde{v},\quad\tilde{v}(t=0)={\bf R}u,\quad t\in[0,\tau],\quad\quad$
$\quad v=(\gamma_{0}+[\gamma_{x},\gamma_{y}]^{\top}{\bm{\nabla}\,})\tilde{v}({\bf x},\tau).$ (16)
Finally, the output $v({\bf x})$ is obtained by taking a directional derivative of $\tilde{v}(\tau,{\bf x})$ in a direction ${\bm{\gamma}}$. All the processes above are encompassed by the mesh-free convolution operator ${\cal C}({\bm{\xi}})$.
The operator ${\cal L}({\bm{\alpha}})={\rm diag}\left({\cal L}({\bm{\alpha}}_{1}),\ldots,{\cal L}({\bm{\alpha}}_{d})\right)$ is an Elliptic Differential Operator, that we define as
@@
The trainable parameters are ${\bm{\xi}}=[\tau,{\bm{\alpha}},{\bm{\gamma}}]\in\mathbb{R}^{7}$, where $\tau$ is a scalar, ${\bm{\gamma}}=[\gamma_{0},\gamma_{x},\gamma_{y}]$, and ${\bm{\alpha}}=[\alpha_{xx},\alpha_{xy},\alpha_{yy}]$. We enforce $\alpha_{xx},\alpha_{yy}>0$ and $\alpha_{xx}\alpha_{yy}-\alpha_{xy}^{2}>0$, ensuring ${\cal L}$ is elliptic and invertible (Trottenberg et al., 2001).
@@
Integrating the PDE over time $[0,\tau]$ and differentiating in the direction $[\gamma_{x},\gamma_{y}]$ effectively applies the convolution to the input $u$. The parameter $\tau$ acts as a scaling factor: increasing it widens and smooths the kernel, while as $\tau\to 0$, the kernel becomes more localized. To visualize our convolution, we sample random parameters ${\bm{\xi}}$ and plot the resulting MFC filters in Figure 3. By varying ${\bm{\xi}}$, we generate a range of convolutional kernels, from local to global, with different orientations. Here, $\alpha$ controls orientation, $\gamma$ sets the lobe direction, and $\tau$ determines the kernel width. The figure demonstrates that our choice of convolutions yields trainable filters to the ones that are typically obtained by CNN (Goodfellow et al., 2016). These filters can change their field of view by changing $\tau$, change their orientation by changing ${\bm{\xi}}$, and change their side lobes’ strength and direction by changing ${\bm{\gamma}}$.
@@
Figure 3: An illustration of Mesh-Free Convolutions (MFCs) to showcase the effect of parameter variations. First row: The parameters are set as $\alpha_{xx}=1$, $\alpha_{yy}=3$, while $\alpha_{xy}$ varies from 0 to 1, and ${\bm{\gamma}}_{x}$ changes from 0 to 1. Second row: The parameters are set as ${\bm{\gamma}}_{x}={\bm{\gamma}}_{y}=0$, ${\bm{\gamma}}_{0}=1$, with $\alpha_{xy}$ varying from 0 to 1, and $t$ changing from 0 to 1. These visualizations highlight the flexibility and expressiveness of MFCs in adapting to different parameter configurations.
@@
Computing MFCs on a discretized mesh. Our MFCs are defined by the solution of Section 3.1 and are continuous functions. Here, we discuss their computation on a discretized mesh for application to images. For a rectangular domain $\Omega$, we use periodic boundary conditions and the Fast Fourier Transform (FFT) to compute the convolution. Let ${\bf u}^{h}$ be a discrete tensor. First, we compute its Fourier transform, $\widehat{\bf u}^{h}={{\bf F}}{\bf u}^{h}$, and use a frequency mesh to obtain the 2D frequency grid ${\bf k}_{x}$ and ${\bf k}_{y}$. The discrete symbol ${\cal L}({\bf k}_{x},{\bf k}_{y})$ is a 3D tensor operating on each channel, similar to standard 2D filters. We then compute a point-wise product and use the inverse FFT to return to image space. Interestingly, given some fixed mesh of size $h$ and a regular 2D convolution ${\bf C}$ it is always possible to build an MFC, ${\bf C}_{{\bm{\xi}}}$, such that they are equivalent when evaluated on that mesh. The proof is given in Appendix B, and it implies that MFCs have at least the same expressiveness as standard convolutions.
@@ 3.2 Fourier analysis of Mesh-Free Convolutions
Following the analysis path shown in Section 2.2, consider the linear loss obtained by these MFCs when discretizing the functions $u$ and $y$. In the case of functions, such as with our MFCs, the inner product in Equation 13 is replaced by an integral,
@@
We study the behavior of our MFC, denoted by ${\cal C}({\bm{\xi}})$, on a function $u(x)$, by computing its symbol. Because the heat equation is linear, the symbol of ${\cal C}({\bm{\xi}})$ is given by (see (Trottenberg et al., 2001)),
${\cal L}(k_{x},$ $k_{y})=\left(\gamma_{0}+ik_{x}\gamma_{x}+ik_{y}\gamma_{y}\right)\cdot$
$\exp\left(-\tau(\alpha_{xx}k_{x}^{2}+2\alpha_{xy}k_{x}k_{y}+\alpha_{yy}k_{y}^{2})\right).$ (19)
@@
Lemma 3.1 (MFCs yield consistent gradients independently of mesh resolution.).
@@
The property described in Lemma 3.1 carries forward to any discrete space, i.e., chosen mesh. Namely, by discretizing our MFCs on a uniform mesh with mesh-size $h$ we obtain its discrete analog:
@@
$\ell^{h}({\bm{\xi}})={\frac{1}{n}}({\bf F}{\bf y}^{h})^{\top}{\cal L}(k_{x},k_{y})({\bf F}{\bf u}^{h})$
$={\frac{1}{n}}\sum_{k_{x},k_{y}}{\cal L}(k_{x},k_{y})\odot\widehat{\bf y}^{h}(k_{x},k_{y})\odot\widehat{\bf u}^{h}(k_{x},k_{y}).$ (21)
As both $\widehat{\bf y}^{h}$ and $\widehat{\bf u}^{h}$ are discrete Fourier transforms of the functions $y$ and $u$, they converge to their continuous counterparts as $h\rightarrow 0$. Thus, $\ell^{h}$ converges to $\ell^{*}$ as $h\rightarrow 0$, leading to a key result: unlike standard convolutions, the gradients of our MFCs converge to a finite value as the mesh refines, making them suitable for multiscale techniques like Multiscale-SGD, summarized in Lemma 3.2.
Lemma 3.2 (MFCs gradients converge as the mesh resolution refines).
Let $\ell^{h}({\bm{\xi}})$ be defined by Equation 20. At the limit,
@@
Table 2: The discrepancy between two consecutive gradients at different resolutions for the loss in Equation 13 for mesh-free convolutions (MFCs). The discrepancy is reduced for both noisy and smooth data.
Computed loss on different resolutions (×$\times 10^{-3}$)
@@
Noisy input 0.16 0.32 0.61 1.3 2.9 5.9
Smooth input 0.11 0.21 0.42 0.78 1.5 3.1
@@ 4 The Full-Multiscale Algorithm
where $C$ is a constant, ${\bm{\theta}}_{*}^{h}$ is the parameter that optimizes the expectation of the loss, and $C_{0}$ is a constant that depends on the initial error at ${\bm{\theta}}_{0}$. Let ${\bm{\theta}}_{*}^{H}$ and ${\bm{\theta}}_{*}^{h}$ be the parameters that minimize the expectation of the loss for meshes with resolution $H>h$ and assume that $\|{\bm{\theta}}_{*}^{H}-{\bm{\theta}}_{*}^{h}\|\leq\gamma H$, where $\gamma$ is a constant. This assumption is justified if the loss converges to a finite value as $h\rightarrow 0$. During the Full-Multiscale-SGD iterations (Algorithm 2), we solve the problem on the coarse mesh $H$ to initialize the fine mesh $h$. Thus, after $k$ steps of SGD on the fine mesh, we can bound the error by,
$\Big{|}{\mathbb{E}}\left[{\ell}({\bm{\theta}}_{k})\right]-{\mathbb{E}}\left[{\ell}({\bm{\theta}}_{*}^{h})\right]\Big{|}\leq\gamma H{\frac{C}{k}}.$ (23)
Requiring that the error is smaller than some $\epsilon$, renders a bounded number of required iterations $k\approx\gamma C{\frac{H}{\epsilon}}$. Since in Algorithm 2, $H=2^{j}h$, the number of iterations for a fixed error $\epsilon$ is halved at each level, the iterations on the finest mesh are a fraction of those on the coarsest mesh. As shown in Table 3, optimal implementation of multiscale can speed up training by a factor of $\sim$160 compared to standard single-scale training.